a person walks a certain distance at a rate of 3.8mph, another person takes 2.1hrs longer to walk the same distance at a rate of 3.2mph. What distance did they walk?
Answer by ankor@dixie-net.com(22740)

a person walks a certain distance at a rate of 3.8mph, another person takes 2.1hrs longer to walk the same distance at a rate of 3.2mph.
What distance did they walk?
:
Let t = time for the faster person to walk the distance
then
(t+2.1) time for the slower person to walk it
:
Write a distance equation: dist = speed * time
they walk the same distance
:
3.8t = 3.2(t+2.1)
3.8t = 3.2t + 6.72
3.8t - 3.2t = 6.72
.6t = 6.72
t = 6.72/.6
t = 11.2 hrs
;
Find the distance using the faster time
11.2 * 3.8 = 42.56 mi
:
Check the distance using the slower time
13.3 * 3.2 = 42.56 mi
: